# Autonomous Issue Resolver: Towards Zero-Touch Code Maintenance

### v2

@@ Abstract
Recent advances in Large Language Models have revolutionized function-level code generation; however, repository-scale Automated Program Repair (APR) remains a significant challenge. Current approaches typically employ a control-centric paradigm, forcing agents to navigate complex directory structures and irrelevant control logic. In this paper, we propose a paradigm shift from the standard Code Property Graphs (CPGs) to the concept of Data Transformation Graph (DTG) that inverts the topology by modeling data states as nodes and functions as edges, enabling agents to trace logic defects through data lineage rather than control flow. We introduce a multi-agent framework leveraging a Dual-Graph Consensus mechanism that reconciles data integrity navigation with control flow logic. Our theoretical analysis and case studies demonstrate that this approach resolves the ”Semantic Trap” inherent in standard RAG systems in modern coding agents. We provide a comprehensive implementation in the form of Autonomous Issue Resolver (AIR), a self-improvement system for zero-touch code maintenance that utilizes neuro-symbolic reasoning and uses the DTG structure for scalable logic repair. Our approach directly addresses the core limitations of current AI code-assistant tools and tackles the critical need for a more robust foundation for our increasingly software-dependent world.
@@ 1 Introduction
In the 21st century, software engineering has evolved from a niche discipline into the digital backbone of the global economy, generating trillions in value and employing hundreds of millions of people worldwide. Today, software is the invisible infrastructure that connects commerce, communication, healthcare, transportation, and nearly every facet of modern life. It has become the omnipresent, global aether of our digital civilization. But unlike a fundamental force of nature, software is created by humans, inheriting our limitations and tendency to make mistakes.
The early days of software developed by smaller teams are long gone. Today’s software products are huge, often comprising millions of lines of code and hundreds of dependencies. No single human, nor even large teams, has a comprehensive understanding of such systems. This is the same as trying to build a skyscraper with bare hands. Current development methodologies attempt to manage this complexity by breaking down the problems, but this merely distributes the potential failures among developers. This is because individual developers, working on isolated components, still introduce bugs, inconsistencies, and architectural drift. The consequences are significant: costly delays, security vulnerabilities, system outages, and a drag on innovation. Industry reports consistently highlight that debugging and maintenance consume a staggering 50-70% of development resources [1]. This ”maintenance tax” prevents organizations from focusing on new feature development, trapping them in a cycle of reactive patching.
The path to truly reliable software lies in systematically reducing the dependence on manual intervention for repetitive, error-prone, and complex analytical tasks. AI-based automation offers the potential for consistency, speed, and depth of analysis that exceed human capabilities. By automating many stages of the software development life-cycle (SDLC), from code generation and testing to debugging and refactoring, we can mitigate human error, enforce best practices, and achieve unprecedented levels of quality assurance.
@@
The core blocker for zero-touch maintenance is not code generation, but systemic reasoning: mapping an issue to a minimal, safe set of edits under repository-wide constraints (APIs, invariants, tests, deploy rules). This demands (i) a repository-scale semantic representation (beyond embeddings alone), and (ii) a risk-aware control policy that decides when to stop, where to edit, and how to validate.
@@
Modern models can synthesize non-trivial patches and reason about code structure, yet this progress has not translated into robust performance on realistic repository-level tasks: on benchmarks such as SWE-bench, resolution rates remain modest, especially for open-source models.
@@
We refer to this gap as the context crisis of repository-scale APR. At small scales, LLMs can be given all relevant code in a single prompt. At repository scale, however, the model must operate under strict context limits while the bug’s cause and manifestation are separated across files, layers, and abstractions. Existing tools address this by improving retrieval and by adding agentic planning, but they leave the underlying representation of the repository largely unchanged: a directory of files, occasionally augmented with control-centric graphs.
@@
We argue that this file-first, control-centered view is misaligned with the causal structure of many real-world bugs and how the development is done in the real world. Human developers rarely debug or write new code by reading files. Instead, they trace the flow of data, following a specific object or variable as it is transformed and routed through the system. We propose to make data-first mental model explicit and to turn it into the primary substrate on which a LLM-based repair agent lives.
Our approach has demonstrated good results on several SWE benchmarks, reaching a resolution rate of 87.1% on SWE-Verified benchmark.
The main contributions of this work are as follows:
  • We introduce the Data-First Transformation Graph (DTG), a data-centric representation of software repositories that performs a semantic compression of Code Property Graphs and aligns with human debugging strategies.
  • We establish connections between the DTG, data-centric optimization frameworks, and Reinforcement Learning, motivating a learning regime in which LLMs reason over graph structure.
  • We propose a novel architecture of the DTG-based repair agent and analyze its performance and accuracy on several benchmarks.
2 Problem Overview and Background
The idea of using AI-based automation in software development is not novel, and there are quite a few products. Tools such as GitHub Copilot, Amazon Code Whisperer, Tabnine [2], Cursor [3], and various chatbot assistants (such as specialized extensions of Claude Code or ChatGPT) are popular today and represent the current state of the art (see also Appendix, Table B.2). Most existing products are iterating on developer help.
@@ 2 Problem Overview and Background
Table B.2 separates professional, basic, and specialized offerings. Professional tools optimize ergonomics for humans-in-the-loop (IDE integration, fast completions), but remain fundamentally reactive and local in scope. Basic tools accelerate scaffolding and prototyping, yet rarely manage repository-scale dependencies or change risk. Specialized systems reach SOTA in narrow settings (e.g., competitive programming or algorithm discovery), but are not designed for end-to-end maintenance workflows.
@@
Table 2.1: Capability taxonomy: where existing tools help - and where they fall short for autonomous maintenance. We classify existing tools by their primary interaction model and identify the specific limitations that prevent them from achieving zero-touch maintenance.
@@ 2.1 Limitations of existing tools
Current LLM-based agents such as SWE-Agent and OpenHands typically interact with repositories through a file-system interface: they list directories, open files, and issue search commands. Each action consumes context and forces the model to maintain a growing internal representation of control and data dependencies. As repositories grow, this misalignment becomes the main bottleneck: the model ”thinks” at the level of dataflow but sees the world as files and control structures.
Retrieval-Augmented Generation (RAG) is the standard strategy for scaling LLMs to large corpora. In code domains, RAG usually combines lexical search with dense embeddings of code snippets. For task prompts involving a particular entity (e.g., fixing some issue that involves a User class), the system retrieves semantically similar snippets for the model to condition on.
@@
  • In data pipelines, failures may be caused by a shared transformation used across multiple domains. Semantically similar code may not share causal influence on the failing data.
This is becoming increasingly common in modern codebases which are often use loose-coupling design patterns when the code tends to live in semantically disconnected clusters of functions which are not called directly (e.g. aspect-oriented programming, dependency injection, and so on).
Traditional RAG treats code as an unordered bag of textual fragments. Even graph-based retrieval (e.g., GraphRAG) typically uses graph structure only to improve retrieval quality, not as the state space in which the agent itself moves. The agent still lives in the file system, reading linearized text.
3 Related Work
The field of LLM-based software engineering has exploded between 2024 and 2026, characterized by a dichotomy between ”Agentic” workflows (complex, multi-step reasoning) and ”Agentless” approaches (simplified, search-and-repair pipelines). The introduction and further development of Transformer-based models trained on large code corpora has revolutionized software engineering. Early breakthroughs like Codex [4] and AlphaCode [5] demonstrated that models could synthesize functional code from natural language descriptions. These foundational models power a generation of interactive assistants, including GitHub Copilot [6], Amazon CodeWhisperer [7], and Cursor [8]. Although effective for local ”next-block” completion [2], these tools primarily operate as ”copilots,” relying on the human developer to manage the broader repository context and verify correctness.
Agentic frameworks such as ReAct [9] and Reflexion [10], allowed LLMs to plan, execute tools, and observe results. State-of-the-art systems like SWE-agent [11] and OpenHands [12] utilize these patterns to interact with repositories via shell commands and file editors, and are currently often used as solid baselines. Similarly, AutoCodeRover [13] employs abstract syntax tree (AST) analysis to improve fault localization.
However, these agents largely treat the repository as a file system to be searched and read, a ”file-first” perspective that suffers from inefficiency when bugs span multiple abstraction layers. Recent evaluations on benchmarks such as SWE-bench [14, 15] indicate that while agents can solve isolated tasks, they struggle with long-horizon reasoning required for enterprise issues [16]. Claude Code [17] and LocAgent [18] attempt to mitigate this by improving retrieval and localization, yet they still fundamentally operate on text-based or partially structured representations that do not explicitly model data lineage.
@@ 3 Related Work
Recent efforts have attempted to integrate such graphs with LLMs to improve repository understanding. CodexGraph [20] and RepoGraph [21], [22] utilize graph databases to enhance Retrieval-Augmented Generation (RAG) by capturing dependencies between files and functions. The Code Graph Model (CGM) [23] similarly integrates graph structures to aid in logical reasoning. However, these approaches often retain the ”representation bloat” of standard CPGs or focus on improving retrieval recall rather than navigation that mimics the ”operational reasoning”. Unlike these control-centric models, our approach utilizes a Data-First Transformation Graph (DTG), which compresses the topology to focus on data states and transformations, somehow aligning our approach with the ”flow-based” mental models used in high-performance computing optimization.
Although most current agents rely on prompt engineering and in-context learning, Reinforcement Learning (RL) has shown promise in discovering novel strategies in complex domains. DeepMind’s work on AlphaZero [24] and subsequently on AlphaDev [25] demonstrated that RL agents could discover algorithms that outperform human-written implementations. In the domain of program repair, early learning-based systems such as Getafix [26] mined fix patterns from static analysis. More recently, AlphaEvolve [27] pushed this further by evolving code structures. Our work incorporates an RL-driven control policy to navigate the DTG, distinguishing AIR from purely prompt-driven agents by allowing it to learn stopping conditions and navigation strategies that minimize regression risks. We are using a simple DQN-based online learning. Note that RL is not a key element of our algorithm, and it can be replaced by simple Monte Carlo Tree Search style graph traversing.
@@ 4.1 A data-first graph representation
Figure 4.1: The partial visualization of the multi-dimensional Data Transformation Graph for a snippet of code. Nodes contains full description of processed data. Edges are effectively transformation functions (sanitize, handle). The arrows demonstrate the ”causal path” of the variable id as it traverses from the process_transactions scope (blue) into the sanitize scope (yellow) and returns as a new entity. This structure achieves Semantic Compression, abstracting away syntactic noise (AST nodes) and enabling the agent to navigate logic defects purely through data lineage.
We propose a fundamental re-architecture of the agent’s environment. Rather than representing a repository as text files, or even as a multi-layer code property graph in which functions and statements are the primary nodes, we focus directly on the data that flows through the system.
We define the Data-First Transformation Graph (DTG) as a directed multigraph with:
@@
To illustrate this concept, consider the following snippet of code that implements a toy transaction handle. Figure 4.1 depict a simplified DTG for this code.
@@
4.2 Why Reinforcement Learning is crucial
The second key ingredient of our design is a smart integration of RL patterns. We draw inspiration from the successes of systems like DeepMind’s AlphaGo and AlphaCode [24][28], which demonstrated how machines can achieve superhuman performance in complex domains not just by mimicking data, but by learning optimal strategies through interaction and self-correction.
@@ 4.2 Why Reinforcement Learning is crucial
4.3 Why holistic reasoning is the key
An RL loop alone is insufficient if the agent cannot understand the code. Generating random code is inefficient, because the size of problem space (e.g. the number of possible variations of code snippet) is bigger than the number of particles in all multiverses combined. We need a powerful and scalable reasoning engine that provides the agent with a deep and structured view of the code base.
@@ 5.1 Code property graphs and representation bloat
The Code Property Graph (CPG), popularized by tools such as Joern [29] and PyXray [30], unifies three classic program representations in a single multi-graph: abstract syntax trees (ASTs), control-flow graphs (CFGs), and program dependence graphs (PDGs). In principle, CPGs contain all information needed for both security analysis and program repair.
@@ 5.4.1 Formal definition
Let a software repository $\mathcal{R}$ be represented as a directed multigraph
$\mathcal{G}_{\mathrm{DTG}}=(V,E),$
where $V$ is a set of vertices (data states - see below) and $E$ is a set of edges (transformations).
@@
  • full_name: A unique identifier, e.g., module.func.var.
  • kind: Class of data, e.g. is it the argument of a function, a global variable, or constant, etc.
  • type: Its static or inferred language-level type, e.g., List[int] or torch.Tensor.
@@ 6 Graph construction pipeline
Constructing the DTG is the most technically demanding phase. We use a ”lazy-loading” pipeline that combines static and dynamic information.
@@
Tree-sitter [31] provides incremental, multi-language AST parsing for languages such as Python, Java, JavaScript, and C++. We use it to:
@@
Custom Tree-sitter queries (in .scm files) define language-specific patterns for these elements [31], but the downstream DTG schema is language-agnostic. Rather than directly exposing the full internal graph of these tools to the LLM, we project only the data-state and transformation information needed for DTG navigation. We use a neo4j graph database as persistent storage and efficient indexing over node and edge attributes.
A key theoretical advantage of the DTG is Semantic Compression. A standard CPG for a function def foo(x): if x>0: return x+1 might contain 20+ nodes (AST nodes for If, Compare, Return, BinOp, etc.). [19] The DTG compresses this entire subgraph into a single edge: $v_{in}\xrightarrow{foo[x>0]}v_{out}$ This reduces the graph size by orders of magnitude (approx. 100x), enabling the agent to load the ”skeleton” of the entire repository into its working memory, solving the context window bottleneck.
Pure data-flow graphs lose control-flow information (branches). We reintroduce this via Guarded Edges. An edge is annotated with a guard $\gamma$ derived from the CFG.
$v_{j}=T(v_{i})\iff\gamma(v_{control})=\text{True}$
This allows the agent to reason: ”This transformation only happens if the is_verified flag is set,” maintaining the necessary logic for bug reproduction without the bloat of the full CFG.

### v3

# WWW home page: https://akaliutau.github.io/

@@ Abstract
Recent advances in Large Language Models have revolutionized function-level code generation; however, repository-scale automated program repair remains difficult because the cause of a defect and its observable failure are often distributed across files, layers, and abstractions, while LLM agents operate under limited context. This paper presents Autonomous Issue Resolver (AIR), a multi-agent repair system centered on a Data-First Transformation Graph (DTG), which represents data states as nodes and transformations as edges. The representation enables the agent to localize faults by following the data line rather than browsing files or relying only on lexical retrieval. AIR combines DTG navigation, targeted editing, and test-driven validation under a risk-aware control policy. The system is evaluated on SWE-bench Verified/Lite and on an exploratory set of 100 issues from 10 open-source Python repositories. AIR achieves 87.1% on SWE-bench Verified and 73.5% on SWE-bench Lite; in the OSS study it resolves 70/100 issues under a fixed tool and a time budget. The ablation results show that the removal of RL reduces the resolution from 87.1% to 82.5%, while the removal of DTG reduces it to 56.2%, indicating that data-first repository representation is the main performance driver. These results suggest that explicit data-lineage navigation is a promising alternative to file-first retrieval for autonomous software maintenance.
keywords:
Automated Program Repair, repository-scale reasoning, data transformation graph, LLM agents, Data Lineage
@@ 1 Introduction
In the 21st century, software engineering has evolved from a niche discipline into the digital backbone of the global economy, generating trillions in value and employing hundreds of millions of people worldwide. Today, software is the invisible infrastructure that connects commerce, communication, healthcare, transportation, and nearly every facet of modern life.
Software maintenance dominates engineering effort: debugging and maintenance account for 50–70% of development resources in many organizations [7]. However, repository-scale automated program repair remains difficult because the cause of a failure and its manifestation are often distributed across files, layers, and abstractions. Current LLM-based coding agents can generate plausible local patches, but remain less reliable when the task requires repository-level fault localization, constrained editing, and regression-aware validation.
@@
The central challenge is not code generation in isolation, but systemic reasoning: mapping an issue report to a minimal, safe set of edits under repository-wide constraints such as APIs, invariants, tests, and deployment rules. This requires both a repository representation that preserves causal structure beyond lexical similarity and a control policy that decides what to inspect, where to edit, and when to stop.
@@
Modern models can synthesize non-trivial patches and reason about code structure, yet this progress has not translated into robust performance on realistic repository-level tasks: on benchmarks such as SWE-bench [8], resolution rates remain modest, especially for open-source models.
@@
This paper refers to this gap as the context crisis of repository-scale APR. At small scales, LLMs can receive all relevant code in a single prompt. At repository scale, however, the model must operate under strict context limits while the bug’s cause and manifestation are separated across files, layers, and abstractions. Existing tools address this by improving retrieval and by adding agentic planning, but they leave the underlying representation of the repository largely unchanged: a directory of files, occasionally augmented with control-centric graphs.
@@
This work argues that this file-first, control-centered view is misaligned with the causal structure of many real-world bugs and how the development is done in the real world. Human developers rarely debug or write new code by reading files. Instead, they trace the flow of data, following a specific object or variable as it is transformed and routed through the system. This paper proposes making the data-first mental model explicit and turning it into the primary substrate on which a LLM-based repair agent lives.
The proposed approach is evaluated on SWE-bench Verified/Lite and an exploratory OSS study; the main results are reported in Section 8.
The main contributions of this work are as follows.
  • A Data-First Transformation Graph (DTG) that models data states as nodes and transformations as edges, enabling causal fault localization beyond file-first retrieval.
  • AIR, a multi-agent repair architecture that combines DTG navigation, targeted editing, and test-driven validation under a risk-aware control policy.
  • An empirical evaluation on the SWE-bench and an exploratory OSS repository study, with isolation of the effects of DTG and the control policy.
The remainder of the paper first summarizes the limits of current coding assistants and APR systems, then positions AIR against related work, introduces the DTG representation and AIR architecture, and finally evaluates the approach on public benchmarks and an exploratory repository-level OSS study, followed by explanations, limitations, and future work.
2 Related Work
2.1 Background and Motivation
AI-assisted software development is already well established in practice. Tools such as GitHub Copilot  [^1] https://docs.github.com/copilot, Amazon CodeWhisperer  [^2] https://docs.aws.amazon.com/codewhisperer/, Tabnine  [^3] https://docs.tabnine.com/main, Claude  [^4] https://claude.ai, and chat-based coding assistants are widely used; however, most of them still optimize developer assistance rather than autonomous repository-scale maintenance (see also Appendix, Table 9). Most existing products are iterating on developer help.
@@ 2.1 Background and Motivation
Table 9 separates professional, basic, and specialized offerings. Professional tools optimize ergonomics for humans-in-the-loop (IDE integration, fast completions), but remain fundamentally reactive and local in scope. Basic tools accelerate scaffolding and prototyping, yet rarely manage repository-scale dependencies or change risk. Specialized systems reach SOTA in narrow settings (e.g., competitive programming or algorithm discovery), but are not designed for end-to-end maintenance workflows.
@@
Table 1: Capability taxonomy: where existing tools help - and where they fall short for autonomous maintenance. We classify existing tools by their primary interaction model and identify the specific limitations that prevent them from achieving zero-touch maintenance.
@@ 2.2 Limitations of existing tools
Current LLM-based agents, such as SWE-Agent and Claude typically interact with repositories through a file-system interface: they list directories, open files, and issue search commands. Each action consumes context and forces the model to maintain a growing internal representation of control and data dependencies. As repositories grow, this misalignment becomes the main bottleneck: the model ”thinks” at the level of dataflow but sees the world as files and control structures.
Retrieval-Augmented Generation (RAG) is the standard strategy for scaling LLMs to large corpora. In code domains, RAG usually combines a lexical search with dense embeddings of code snippets. For task prompts involving a particular entity (e.g., fixing some issue that involves a User class), the system retrieves semantically similar snippets for the model to condition on.
@@
  • In data pipelines, failures may be caused by a shared transformation used in multiple domains. Semantically similar code snippets may not share a causal influence on the failing data.
This is increasingly common in modern codebases that use loosely coupled design patterns, where relevant logic may reside in semantically disconnected function clusters that are not called directly (e.g. aspect-oriented programming, dependency injection, and so on).
Traditional RAG treats the code as an unordered bag of textual fragments. Even graph-based retrieval (e.g., GraphRAG) typically uses graph structure only to improve retrieval quality, not as the state space in which the agent itself moves. The agent still lives in the file system, reading linearized text.
Taken together, prior work improves code generation, retrieval, or tool use, but the dominant interaction model remains file-first: graphs are typically used to improve retrieval, while the agent still plans and edits in a text/file workspace. AIR differs in two respects. First, the DTG is the agent’s primary navigation space, rather than an auxiliary retrieval index. Second, the control policy is tied to validation outcomes, with the goal of reducing over-editing and premature patch acceptance. This positions AIR as a repository-scale repair system focused on causal localization and change risk, rather than as a general-purpose coding assistant.
2.3 Literature Review
Transformer-based models trained on large code corpora have substantially advanced software engineering. Early systems such as Codex [6] and AlphaCode [11] showed that models can synthesize functional code from natural-language descriptions. These advances underpin modern interactive coding assistants. However, while such systems are effective for local ”next-block” completion [23], they still rely on the human developer to manage repository-wide context and verify correctness.
Agentic frameworks such as ReAct [22] and Reflexion [15], allowed LLMs to plan, execute tools, and observe outputs. State-of-the-art systems like SWE-agent [21] and OpenHands [19] utilize these patterns to interact with repositories via shell commands and file editors, and currently often used as solid baselines. Similarly, AutoCodeRover [25] employs abstract syntax tree (AST) analysis to improve fault localization.
However, these agents largely treat the repository as a file system to be searched and read, a file-first perspective that becomes inefficient when defects span multiple abstraction layers. Evaluations on SWE-bench indicate that such agents can solve isolated tasks but remain weaker on long-horizon repository-level reasoning [8, 24]. More recent localization-oriented systems, such as LocAgent [4], improve retrieval and fault localization, yet they still operate primarily on text-based or partially structured representations rather than on explicit data lineage.
@@ 2.3 Literature Review
Recent efforts have attempted to integrate such graphs with LLMs to improve repository understanding. CodexGraph [3] and RepoGraph [2] utilize graph databases to enhance Retrieval-Augmented Generation (RAG) by capturing dependencies between files and functions. The Code Graph Model (CGM) [1] similarly integrates graph structures to aid in logical reasoning. However, these approaches often retain the representation bloat of standard CPGs or focus on improving retrieval recall rather than navigation. Unlike these control-centric models, the proposed approach utilizes a Data-First Transformation Graph (DTG), which compresses the topology to focus on data states and transformations and is aligned with flow-based debugging and optimization practice.
Although most current agents rely on prompt engineering and in-context learning, Reinforcement Learning (RL) has shown promise in discovering novel strategies in complex domains. DeepMind’s work on AlphaZero [16] and AlphaDev [12] demonstrated that RL agents can discover algorithms that outperform human-written implementations. In the domain of program repair, early learning-based systems such as Getafix [5] mined fix patterns from static analysis. AIR incorporates an RL-driven control policy to navigate the DTG, allowing the system to learn stopping conditions and navigation strategies that minimize regression risks. A simple DQN-based online learning scheme is used; RL is not a core dependency or main focus of the proposed method and could be replaced for example by simpler Monte Carlo Tree Search-style graph traversal.
@@ 3.1 A data-first graph representation
Figure 1: The partial visualization of the multi-dimensional Data Transformation Graph for a snippet of code. Nodes contains full description of processed data. Edges are effectively transformation functions (sanitize, handle)
This paper proposes a fundamental re-architecture of the agent’s environment. Rather than representing a repository as text files, or even as a multi-layer code property graph in which functions and statements are the primary nodes, we focus directly on the data that flows through the system.
The Data-First Transformation Graph (DTG) is defined as a directed multigraph with:
@@
To illustrate this concept, consider the following snippet of code that implements a toy transaction handle. Figure 1 depict a simplified DTG for this code.
@@
3.2 Role of the control policy
In AIR, reinforcement learning is used to learn navigation and stopping decisions, not to generate code tokens directly. Given the current DTG neighborhood, the policy selects actions such as navigating upstream or downstream, inspecting a transformation, invoking validation, or terminating the search. The rewards are tied to successful issue resolution, regression avoidance, and editing efficiency. This makes the control policy risk-aware: trajectories that produce broad or unstable patches are penalized relative to minimal validated fixes.
The second key ingredient is the integration of RL patterns. The design draws inspiration from the successes of systems like DeepMind’s AlphaGo and AlphaCode [16][10], which demonstrated how machines can achieve superhuman performance in complex domains not just by mimicking data, but by learning optimal strategies through interaction and self-correction.
@@ 3.2 Role of the control policy
3.3 Why structured repository reasoning is necessary
A learned control policy alone does not solve repository-scale repair because the search space remains combinatorial. AIR therefore couples the policy with DTG-based context. The graph narrows exploration to causally relevant data lineages, while the policy determines which neighborhoods to inspect and when enough evidence has been accumulated to move from localization to editing and validation. This division of labor is what makes the approach scalable at repository level.
@@ 4.1 Code property graphs and representation bloat
The Code Property Graph (CPG), popularized by tools such as Joern [9], unifies three classic program representations in a single multi-graph: abstract syntax trees (ASTs), control-flow graphs (CFGs), and program dependence graphs (PDGs). In principle, CPGs contain all information needed for both security analysis and program repair.
@@ 4.4.1 Formal definition
Let a software repository $\mathcal{R}$ be represented as a directed multigraph G_DTG= (V, E), where $V$ is a set of vertices (data states - see below) and $E$ is a set of edges (transformations).
@@
  • name: A unique identifier, e.g., module.func.var.
  • kind: Class of data, e.g. is it the argument of function, a global variable, or constant, etc.
  • type: Its static or inferred language-level type, for example, List[int] or torch.Tensor.
@@ 5 Graph construction pipeline
Constructing the DTG is the most technically demanding phase. The current implementation uses a ”lazy-loading” pipeline that combines static and dynamic information.
@@
Tree-sitter [18] provides incremental, multi-language AST parsing for languages such as Python, Java, JavaScript, and C++. It is used to:
@@
Custom Tree-sitter queries (in .scm files) define language-specific patterns for these elements [18], but the downstream DTG schema is language-agnostic. Rather than directly exposing the full internal graph of these tools to the LLM, we project only the data-state and transformation information needed for DTG navigation. A neo4j graph database is used as persistent storage and efficient indexing over node and edge attributes.